OpenStudy (anonymous):

A system of equations is shown below: 6x = -7y + 11 5x + 2y = 13 What is the solution to this system of equations? (-3, 1) (-1, 3) (1, -3) (3, -1)

OpenStudy (anonymous):

Some help would be amazing

OpenStudy (anonymous):

im looking at it hold on

OpenStudy (anonymous):

thank you

OpenStudy (uri):

Arrange the first equation like x+y=(number)

OpenStudy (uri):

6x = -7y + 11 6x+7y=11 <-

OpenStudy (uri):

Then you have to find x and y.

OpenStudy (anonymous):

its (3,-1)

OpenStudy (uri):

@epicjackal117 Please don't give direct answers. We are here to help.

OpenStudy (anonymous):

well this is the easiest way, start plugging stuff in

OpenStudy (anonymous):

cuz look 6*3=18 and 7*-1=-7 so it leaves you with 11

OpenStudy (anonymous):

and when uri showed you about moving the equation to the other side, that 11 turns negative

OpenStudy (anonymous):

(3,-1) is the correct solution of the above equations i.e x =3 and y -1

OpenStudy (anonymous):

wait never mind sorry

OpenStudy (anonymous):

and 18-7=11 so its (3,-1)

OpenStudy (anonymous):

uri was correct

OpenStudy (anonymous):

sorry guys my computer went crazy

OpenStudy (anonymous):

once both equations are properly set up you should be able to subtract like terms

OpenStudy (anonymous):

and end up with one equation

OpenStudy (anonymous):

once you end up with the said equation, you plug in the coordinate points and it both sides of the equation should be equal.

OpenStudy (anonymous):

thank you so so much!!

OpenStudy (anonymous):

no problem, happy to help!

OpenStudy (anonymous):

np